# thanks: Corresponding authorthanks: Corresponding author

### v1

@@ Abstract
Quantum steering is a crucial quantum resource that lies intermediate between entanglement and Bell nonlocality. Gaussian channels, meanwhile, play a foundational role in diverse quantum protocols, secure communication, and related fields. In this paper, we focus on several classes of Gaussian channels associated with quantum steering: Gaussian steering-annihilating channels, Gaussian steering-breaking channels, Gaussian unsteerable channels, and maximal Gaussian unsteerable channels. We give the concepts of these channels, derive the necessary and sufficient conditions for a Gaussian channel to belong to each class, and explore the intrinsic relationships among them. Additionally, since quantifying the steering capability of Gaussian channels in continuous-variable systems requires an understanding of the structure of free superchannels, we also provide a detailed characterization of Gaussian unsteerable superchannels and maximal Gaussian unsteerable superchannels.
@@ I Introduction
Gaussian states are a special class of quantum states in continuous-variable (CV) systems, playing a pivotal role in quantum optics and quantum information theory (SLP ; WPG1 ; S1 ). Over the past few years, the EPR steering criteria and measures for Gaussian states have garnered considerable attention from researchers (see, e.g., JiL ; KSA ; KLRA ; ZYL ; TGRF ; MKIB ; BJC ; YHQ and the references therein). Notably, quantum information processing inevitably involves quantum channels. As a distinctive class of quantum channels, Gaussian quantum channels not only furnish a core theoretical framework for elucidating the inherent physical limitations of quantum communication and quantum computing, but also directly underpin the translation of quantum technologies from theoretical concepts to practical applications (BSS ; MDG ; CEGH ; CD ; NGA ; LD ; PMGH ). They thus hold irreplaceable significance in quantum systems, especially in optical quantum systems. One primary objective of this paper is to investigate three types of special Gaussian quantum channels with respect to EPR steering: Gaussian steering-annihilating channels that completely eliminate steering; Gaussian steering-breaking channels that locally disrupt steering; and maximal Gaussian unsteerable channels that map Gaussian unsteerable states into Gaussian unsteerable states.
@@ II.1 Gaussian states
is called the mean or the displacement vector of $\rho$ and $\Gamma=(\gamma_{kl})\in\mathcal{M}_{2N}(\mathbb{R})$ is called the covariance matrix (CM) of $\rho$ defined by $\gamma_{kl}={\rm tr}[\rho(\Delta\hat{R}_{k}\Delta\hat{R}_{l}+\Delta\hat{R}_{l}\Delta\hat{R}_{k})]$ with $\Delta\hat{R}_{k}=\hat{R}_{k}-\langle\hat{R}_{k}\rangle$ (SLP ). Here, $\mathcal{M}_{d}(\mathbb{R})$ stands for the algebra of all $d\times d$ matrices over the real field $\mathbb{R}$. So, any Gaussian state $\rho$ with CM $\Gamma$ and displacement vector $\mathbf{d}$} will sometimes be represented as $\rho(\Gamma,{\mathbf{d}})$. Note that $\Gamma$ is real symmetric and satisfies the condition
$\Gamma+i\Omega_{N}\geq 0,\ \ {\rm where}\ \ \Omega_{N}=\oplus_{k=1}^{N}\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$.
@@
In the rest of this paper, if there is no explanation, we always assume that $H$, $H_{A}$ and $H_{B}$ are infinite-dimensional separable complex Hilbert spaces with $H=H_{A}\otimes H_{B}$.
@@ II.2 EPR steering
A measurement assemblage ${\mathcal{MA}}=\{M_{a|x}\}_{a,x}$ is a collection of positive operators $M_{a|x}\geq 0$ satisfying $\sum_{a}M_{a|x}=I$ for each $x$. Such a collection represents one positive-operator-valued measurement (POVM), describing a general quantum measurement, for each $x$. In a (bipartite) EPR steering scenario, one party performs measurements on a shared state $\rho_{AB}\in{\mathcal{S}}(H_{A}\otimes H_{B})$, which steers the quantum state of the other particle. If Alice performs a set of measurements $\{M^{A}_{a|x}\}_{a,x}$, then the collection of sub-normalized of Bob is an assemblage $\{\rho^{B}_{a|x}\}_{a,x}$ with
@@
If every assemblage on Bob $\{\rho^{B}_{a|x}\}_{a,x}$ can be explained by a local hidden state (LHS) model:
@@
where $\lambda$ is a hidden variable, distributed according to $p_{\lambda}$, $\sigma_{\lambda}$ are “hidden states” of Bob, and $p(a|x,\lambda)$ are local “response functions” of Alice, then we say that $\rho_{AB}$ has a LHS form, or does not demonstrate steering (WJ ). Otherwise, if there exist measurements such that $\rho^{B}_{a|x}$ does not admit such LHS decomposition, then $\rho_{AB}$ is called steerable from A to B. Symmetrically, we can define the steerability of $\rho_{AB}$ from B to A. Here, we point out that, if there is no special illustration, we focus on the steerability from A to B in this paper.
@@
where $\alpha=(\alpha_{1},\cdots,\alpha_{N})^{\rm T}\in{\mathbb{C}^{N}}$, $D(\alpha)={\rm exp}[\sum_{j=1}^{N}(\alpha_{j}\hat{a}_{j}^{\dagger}-\alpha_{j}^{*}\hat{a}_{j})]$ is the $N$-mode Weyl displacement operator and $\varpi$ is a zero mean $N$-mode Gaussian state, which is called the seed state of $\Pi$ (GC ).
For any bipartite Gaussian state, the authors in WJ derived a linear matrix inequality steering criterion via GPOVMs. Assume that $\rho\in{\mathcal{S}}(H_{A}\otimes H_{B})$ is any $(m+n)$-mode Gaussian state with CM $\Gamma_{\rho}$ in Eq.(3). As demonstrated in WJ , $\rho$ is unsteerable by the subsystem A’s all GPOVMs if and only if
$\Gamma_{\rho}+0_{2m}\oplus i\Omega_{n}\geq 0,\ \ {\rm where}\ \ \Omega_{n}=\oplus_{k=1}^{n}\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$. (5)
Obviously,
$\Gamma_{\rho}+0_{2m}\oplus i\Omega_{n}\geq 0\Leftrightarrow\Gamma_{\rho}\geq\pm i(0_{2m}\oplus\Omega_{n}).$
Denote respectively by ${\mathcal{GS}}(H)$ and ${\mathcal{GS}}_{\mathcal{US}}(H_{A}\otimes H_{B})$ the set of all Gaussian states on $H$ and the set of all Gaussian unsteerable states from A to B on $H_{A}\otimes H_{B}$, that is,
$\mathcal{GS}(H)=\{\rho\in{\mathcal{S}}(H):\ \rho\ \mbox{is a Gaussian state}\}$
@@
$\begin{array}[]{rl}&{\mathcal{GS}}_{\mathcal{US}}(H_{A}\otimes H_{B})\\ =&\{\rho\in{\mathcal{GS}}(H_{A}\otimes H_{B}):\Gamma_{\rho}+0_{2m}\oplus i\Omega_{n}\geq 0\}.\end{array}$
@@ II.3 Gaussian unsteerable channels
Recall that a Gaussian channel is a quantum channel which transforms Gaussian states into Gaussian states (PMGH ; CEGH ). An $N$-mode Gaussian channel $\phi$ on $H$ can be described by $\phi=\phi\left({K,M,\mathbf{d}}\right)$, which acts on $\rho(\Gamma_{\rho},\mathbf{d}_{\rho})\in\mathcal{GS}(H)$ as
@@
$\mathbf{d}\in\mathbb{R}^{2N}$ is a column displacement vector, $K,M\in\mathcal{M}_{2N}(\mathbb{R})$ satisfy $M=M^{\rm T}$ and the completely positive condition
@@
Denote by ${\mathcal{GC}}(H)$ the set of all Gaussian channels on $H$. Here, we give three known Gaussian channels which are used frequently.
Attenuator channels. A single-mode attenuator channel $\phi_{\theta}^{n_{th}}(K,M,\mathbf{d})$ is a deterministic Gaussian channel with
@@
where $\theta\in[0,2\pi]$ and thermal noise $n_{th}\geq 1$. Particularly, if $n_{th}=1$, then the channel is called a pure lossy channel.
Constant channels. A Gaussian channel $\Theta(K,M,\mathbf{d})\in{\mathcal{GC}}(H)$ is called a constant channel if there exists some $\rho_{0}(\Gamma_{0},{\mathbf{d}}_{0})\in\mathcal{GS}(H)$ such that $\Theta(\rho)=\rho_{0}$ for all $\rho\in\mathcal{GS}(H)$. In this case, $\Theta(K,M,\mathbf{d})$ can be represented as $\Theta(K,M,\mathbf{d})=\Theta(0,\Gamma_{0},\mathbf{\mathbf{d}}_{0})$.
Identity Gaussian channel. A Gaussian channel $\psi(K,M,\mathbf{d})\in{\mathcal{GC}}(H)$ is called an identity channel if $\psi(\rho)=\rho$ holds for all $\rho\in\mathcal{GS}(H)$. In this case, $K=I$, $M=0$ and $\mathbf{d}=0$.
In YHQ , the authors gave the definition of Gaussian unsteerable channels. Any $(m+n)$-mode Gaussian channel $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}(H_{A}\otimes H_{B})$ is called Gaussian unsteerable (from A to B) if $K$ and $M$ satisfy the following relation
@@
and is called maximal Gaussian unsteerable (from A to B) if $\phi({\mathcal{GS}}_{\mathcal{US}}(H_{A}\otimes H_{B}))\subseteq{\mathcal{GS}}_{\mathcal{US}}(H_{A}\otimes H_{B})$. Let
$\begin{array}[]{rl}&{\mathcal{GC}}_{\mathcal{US}}(H_{A}\otimes H_{B})\\ =&\{\phi(K,M,\mathbf{d})\in{\mathcal{GC}}(H_{A}\otimes H_{B}):\\ &M+(0_{2m}\oplus i{\Omega_{n}})-K\left(0_{2m}\oplus i{\Omega_{n}}\right){K^{\rm T}}\geq 0\}\end{array}$
@@
$\begin{array}[]{rl}&{\mathcal{GC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B})\\ =&\{\phi(K,M,\mathbf{d})\in{\mathcal{GC}}(H_{A}\otimes H_{B}):\\ &\phi({\mathcal{GS}}_{\mathcal{US}}(H_{A}\otimes H_{B}))\subseteq{\mathcal{GS}}_{\mathcal{US}}(H_{A}\otimes H_{B})\}.\end{array}$
It is shown in YHQ that the set ${\mathcal{GC}}_{\mathcal{US}}(H_{A}\otimes H_{B})$ is a proper (but) large subset of ${\mathcal{GC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B})$. Obviously, the $(m+n)$-mode identity Gaussian channel $\phi(I,0,0)\in{\mathcal{GC}}_{\mathcal{US}}(H_{A}\otimes H_{B})$.

### v2

@@ Abstract
EPR steering is a crucial quantum resource that lies intermediate between entanglement and Bell nonlocality. Gaussian channels, meanwhile, play a foundational role in diverse quantum protocols, secure communication, and related fields. In this paper, we focus on several classes of Gaussian channels associated with EPR steering: Gaussian steering-annihilating channels, Gaussian steering-breaking channels, Gaussian unsteerable channels, and maximal Gaussian unsteerable channels. We give the concepts of these channels, derive the necessary and sufficient conditions for a Gaussian channel to belong to each class, and explore the intrinsic relationships among them. Additionally, since quantifying the steering capability of Gaussian channels in continuous-variable systems requires an understanding of the structure of free superchannels, we also provide a detailed characterization of Gaussian unsteerable superchannels and maximal Gaussian unsteerable superchannels.
@@ I Introduction
Gaussian states are a special class of quantum states in continuous-variable (CV) quantum systems, playing a pivotal role in quantum optics and quantum information theory SLP ; WPG1 ; S1 . Over the past few years, the EPR steering criteria and measures for Gaussian states have garnered considerable attention from researchers (see JiL ; KSA ; KLRA ; ZYL ; TGRF ; MKIB ; BJC ; YHQ and the references therein). Notably, quantum information processing inevitably involves quantum channels. As a distinctive class of quantum channels, Gaussian quantum channels not only furnish a core theoretical framework for elucidating the inherent physical limitations of quantum communication and quantum computing, but also directly underpin the translation of quantum technologies from theoretical concepts to practical applications BSS ; MDG ; CEGH ; CD ; NGA ; LD ; PMGH . They thus hold irreplaceable significance in quantum systems, especially in optical quantum systems.
In quantum information processing, the storage and transmission of quantum states are crucial but inevitably affected by the environment. Thus, researching which types of environmental influences should be avoided and which are acceptable is of critical importance. In other words, analyzing the structure of quantum channels is essential. One promising research direction focuses on the dynamics of quantum resources, such as quantum entanglement, EPR steering and non-locality, under local quantum channels. Horodecki, Shor and Ruskai HSR studied entanglement-breaking channels in bipartite discrete-variable quantum systems which completely disentangle the subsystem they act on from the rest of the system, and proved that the Holevo capacity of entanglement-breaking channels is additive. This work provides theoretical support for optimizing classical information transmission over noisy channels in quantum communication. Holevo HEB discussed the condition any mode Gaussian quantum channels becoming entanglement-breaking, and using this condition to study several additivity conjectures of capacity for one mode Gaussian channels. The results in KGEB indicate that Gaussian entanglement-breaking channels and entanglement-annihilating channels serve as essential benchmarks for the infeasibility assessment of entanglement preservation, and act as core tools for both theoretical analysis and practical guidance in this field. Based on Gaussian entanglement-breaking channels, passive eavesdropping-immune CV quantum key distribution protocols can be constructed. Pereg P11 investigated the communication performance of entanglement-breaking channels with unreliable entanglement assistance, revealing the mechanism by which initial entanglement enhances the signal-to-noise ratio through classical correlations. Hence the study on (Gaussian) entanglement-breaking channels is of great practical significance.
Note that EPR steering is another important quantum resource. Inspired by the above works, the primary objective of this paper is to investigate some types of special Gaussian quantum channels with respect to EPR steering: Gaussian steering-breaking channels that locally disrupt steering; Gaussian steering-annihilating channels that completely eliminate steering; and maximal Gaussian unsteerable channels that map Gaussian unsteerable states into Gaussian unsteerable states, so as to exploit the ability of Gaussian channels to create or destroy steering and  lay a solid theoretical foundation for the resource theory of steering for Gaussian channels.
@@ II.1 Gaussian states
is called the mean or the displacement vector of $\rho$ and $\Gamma=(\gamma_{kl})\in\mathcal{M}_{2N}(\mathbb{R})$ is called the covariance matrix (CM) of $\rho$ defined by $\gamma_{kl}={\rm tr}[\rho(\Delta\hat{R}_{k}\Delta\hat{R}_{l}+\Delta\hat{R}_{l}\Delta\hat{R}_{k})]$ with $\Delta\hat{R}_{k}=\hat{R}_{k}-\langle\hat{R}_{k}\rangle$ SLP . Here, $\mathcal{M}_{d}(\mathbb{R})$ stands for the algebra of all $d\times d$ matrices over the real field $\mathbb{R}$. So, any Gaussian state $\rho$ with CM $\Gamma$ and displacement vector $\mathbf{d}$} will sometimes be represented as $\rho(\Gamma,{\mathbf{d}})$. Note that $\Gamma$ is real symmetric and satisfies the condition
$\Gamma+i\Omega_{N}\geq 0,$
where
$\Omega_{N}=\underbrace{\Omega\oplus\cdots\oplus\Omega}_{N}\ \ {\rm with}\ \ \Omega=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$.
@@ II.2 EPR steering
In a bipartite EPR steering scenario, Alice and Bob share a bipartite state $\rho_{AB}\in{\mathcal{S}}(H_{A}\otimes H_{B})$, and Alice performs positive-operator-valued measurements (POVMs) on her subsystem to steer $\rho_{AB}$ on Bob’s side. If Alice performs a set of POVMs ${\mathcal{MA}}=\{M_{a|x}\}_{a,x}$ (that is, $M_{a|x}\geq 0$ with $\sum_{a}M_{a|x}=I$ for each $x$), then the assemblage of sub-normalized “conditional states” of the subsystem state $\rho_{B}={\rm Tr}_{A}(\rho_{AB})$ is $\{\rho^{B}_{a|x}\}_{a,x}$, where
@@
$\rho_{AB}$ is said to be unsteerable from A to B if every assemblage $\{\rho^{B}_{a|x}\}_{a,x}$ on Bob’s side can be explained by a local hidden state (LHS) model as follows
@@
where $\lambda$ is a hidden variable, $p_{\lambda}$ is a distribution in $\lambda$, $p(a|x,\lambda)$ are local “response functions” of Alice, and $\sigma_{\lambda}$ are “hidden states” of Bob. Otherwise, $\rho_{AB}$ is called steerable (from A to B) WJ . Symmetrically, we can define the steerability of $\rho_{AB}$ (from B to A).
@@
where $\alpha=(\alpha_{1},\cdots,\alpha_{N})^{\rm T}\in{\mathbb{C}^{N}}$, $D(\alpha)={\rm exp}[\sum_{j=1}^{N}(\alpha_{j}\hat{a}_{j}^{\dagger}-\alpha_{j}^{*}\hat{a}_{j})]$ is the $N$-mode Weyl displacement operator and $\varpi$ is a zero mean $N$-mode Gaussian state, which is called the seed state of $\Pi$ GC .
For any bipartite Gaussian state, the authors in WJ derived a linear matrix inequality steering criterion via GPOVMs. Assume that $\rho\in{\mathcal{S}}(H_{A}\otimes H_{B})$ is any $(m+n)$-mode Gaussian state with CM $\Gamma_{\rho}$ in Eq.(3). As demonstrated in WJ , $\rho$ is unsteerable (from A to B)by the subsystem A’s all GPOVMs if and only if
$\Gamma_{\rho}+0_{2m}\oplus i\Omega_{n}\geq 0;$ (5)
and $\rho$ is unsteerable by the subsystem B’s all GPOVMs if and only if
$\Gamma_{\rho}+i\Omega_{m}\oplus 0_{2n}\geq 0.$
Note that any positive semidefnite matrix and its transpose matrix have the same eigenvalues. So the condition (5) is equivalent to
$(\Gamma_{\rho}+0_{2m}\oplus i\Omega_{n})^{\rm T}\geq 0,$
that is,
$\Gamma_{\rho}-0_{2m}\oplus i\Omega_{n}\geq 0.$
Hence
$\begin{array}[]{rl}&\Gamma_{\rho}+0_{2m}\oplus i\Omega_{n}\geq 0\\ \Leftrightarrow&\Gamma_{\rho}\geq\pm(0_{2m}\oplus i\Omega_{n})=\pm i(0_{2m}\oplus\Omega_{n}).\end{array}$
Denote respectively by ${\mathcal{GS}}^{(N)}$ and $\mathcal{GS}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$} the set of all $N$-mode Gaussian states, the set of all $(m+n)$-mode Gaussian states that are Gaussian unsteerable (from A to B) by using Alice’s Gaussian measurements. Particularly, if $N=m+n$, denote ${\mathcal{GS}}^{(N)}={\mathcal{GS}}^{(m,n)}.$ That is,
${\mathcal{GS}}^{(N)}=\{\mbox{all N-mode Gaussian states}\}$
@@
$\begin{array}[]{rl}&{\mathcal{GS}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}}\\ =&\{\rho\in{\mathcal{GS}}^{(m,n)}:\Gamma_{\rho}+0_{2m}\oplus i\Omega_{n}\geq 0\}\\ =&\{\rho\in{\mathcal{GS}}^{(m,n)}:\Gamma_{\rho}\geq\pm i(0_{2m}\oplus\Omega_{n})\}.\end{array}$
@@ II.3 Gaussian unsteerable channels
Recall that a Gaussian channel is a quantum channel which transforms any Gaussian states into Gaussian statesPMGH ; CEGH . An $N$-mode Gaussian channel $\phi$ can be described by $\phi=\phi\left({K,M,\mathbf{d}}\right)$, which acts on $\rho(\Gamma_{\rho},\mathbf{d}_{\rho})\in\mathcal{GS}(N)$ as
@@
where $\mathbf{d}\in\mathbb{R}^{2N}$ is a column displacement vector, $K,M\in\mathcal{M}_{2N}(\mathbb{R})$ satisfy $M=M^{\rm T}$ and the completely positive condition
@@
Denote by ${\mathcal{GC}}^{(N)}$ the set of all $N$-mode Gaussian channels:
${\mathcal{GC}}^{(N)}=\{\mbox{all N-mode Gaussian quantum channels}\}.$
Particularly, if $N=m+n$, write ${\mathcal{GC}}^{(N)}={\mathcal{GC}}^{(m,n)}$.
Here, we give three known Gaussian channels which are used frequently.
Attenuator channels. A single-mode attenuator channel $\phi_{\theta}^{n_{th}}(K,M,\mathbf{d})$ is a deterministic Gaussian channel S1 with
@@
where $\theta\in[0,2\pi]$, the thermal noise $n_{th}=2N_{0}+1\geq 1$ and $N_{0}$ is mean number of thermal occupation. Particularly, if $n_{th}=1$, then the channel is called a pure lossy channel.
Constant channels. An $N$-mode Gaussian channel $\Theta(K,M,\mathbf{d})$ is called a constant channel if there exists some $N$-mode Gaussian state $\rho_{0}(\Gamma_{0},{\mathbf{d}}_{0})\in\mathcal{GS}^{(N)}$ such that $\Theta(\rho)=\rho_{0}$ for all $N$-mode Gaussian states $\rho\in\mathcal{GS}^{(N)}$. In this case, $\Theta(K,M,\mathbf{d})$ can be represented as $\Theta(K,M,\mathbf{d})=\Theta(0,\Gamma_{0},\mathbf{\mathbf{d}}_{0})$.
Identity Gaussian channel. An $N$-mode Gaussian channel $\psi(K,M,\mathbf{d})$ is called an identity channel if $\psi(\rho)=\rho$ holds for all $N$-mode Gaussian states $\rho\in\mathcal{GS}^{(N)}$. In this case, $K=I$, $M=0$ and $\mathbf{d}=0$.
In YHQ , the authors gave the definition of Gaussian unsteerable channels. Recall that any $(m+n)$-mode Gaussian channel $\phi=\phi\left({K,M,\mathbf{d}}\right)$ is called Gaussian unsteerable from A to B if $K$ and $M$ satisfy the following relation
@@
and is called maximal Gaussian unsteerable from A to B if $\phi$ maps all $(m+n)$-mode Gaussian unsteerable states from A to B into $(m+n)$-mode Gaussian unsteerable states from A to B.
Let ${\mathcal{GC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$ and ${\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}$ stand for the set of all $(m+n)$-mode Gaussian unsteerable (from A to B) channels and the set of all $(m+n)$-mode maximal Gaussian unsteerable (from A to B) channels, respectively, that is,
${\mathcal{GC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}=\{\phi(K,M,\mathbf{d})\in{\mathcal{GC}}^{(m,n)}:$
$M+(0_{2m}\oplus i{\Omega_{n}})-K\left(0_{2m}\oplus i{\Omega_{n}}\right){K^{\rm T}}\geq 0\},$
@@
${\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}=\{\phi(K,M,\mathbf{d})\in{\mathcal{GC}}^{(m,n)}:$
$\phi(\mathcal{GS}^{(m,n)}_{\mathcal{US}(A\rightarrow B)})\subseteq\mathcal{GS}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}\}.$
It is shown in YHQ that the set ${\mathcal{GC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$is a proper (but) large subset of ${\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}$. Obviously, the $(m+n)$-mode identity Gaussian channel $\phi(I,0,0)\in{\mathcal{GC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$.